OpenStudy (anonymous):

Can someone help me simplify this trig expression???............sec(x)cot(x)-cot(x)cos(x)

OpenStudy (unklerhaukus):

sec cot-cot cos sub in sec = 1/cos, and cot = 1/tan = cos/sin

OpenStudy (anonymous):

yeah i did so but its confusing me

OpenStudy (unklerhaukus):

sec cot - cot cos = (1/cos)(cos/sin) - (cos/sin)cos now simplify =

OpenStudy (anonymous):

(1/cosx)(cosx/sinx)-(cosx/sinx * cosx)

OpenStudy (unklerhaukus):

right, now simplify

OpenStudy (anonymous):

yeah thats where im stumped

OpenStudy (unklerhaukus):

the first term is (1/cosx)(cosx/sinx) cancel the cosines

OpenStudy (anonymous):

1/sinx ?

OpenStudy (unklerhaukus):

yeah, now simplify (cosx/sinx * cosx)

OpenStudy (anonymous):

cos^2x/sinx ?

OpenStudy (unklerhaukus):

good, so you have 1/sinx - cos^2x/sinx ? now factor out the 1/sin

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

1-cos^2x/sinx

OpenStudy (anonymous):

so it's sinx?

OpenStudy (unklerhaukus):

yeah (1-cos^2x)/sinx now use sin^2+cos^2=1 i.e. 1-cos^2 = sin^2

OpenStudy (unklerhaukus):

yes!

OpenStudy (anonymous):

cool. one more??

OpenStudy (unklerhaukus):

ok

OpenStudy (anonymous):

cosx/1+sinx + cosx/1-sinx

OpenStudy (unklerhaukus):

is that meant to be cosx/(1+sinx) + cosx/(1-sinx) ?

OpenStudy (anonymous):

yeah

OpenStudy (unklerhaukus):

change the first term by multiplying by 1 = (1-sinx)/(1-sinx) cosx/(1+sinx) = cosx/(1+sinx) * (1-sinx)/(1-sinx)

OpenStudy (anonymous):

i dont know where to start

OpenStudy (unklerhaukus):

cosx/(1+sinx) = cosx/(1+sinx) * (1-sinx)/(1-sinx) = cosx(1-sinx) / (1-sin^2x)

OpenStudy (unklerhaukus):

now do something similar to the second term: multiply cosx/(1-sinx) by 1 = (1+sinx)/(1+sinx)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

cos(1-sinx)/1-sin^2x

OpenStudy (unklerhaukus):

not quite cos(1+sinx)/(1-sin^2x)

OpenStudy (anonymous):

oh my bad thats what i had

OpenStudy (anonymous):

i got 2cosx/cos^2x

OpenStudy (unklerhaukus):

so you have cosx/(1+sinx) + cosx/(1-sinx) = cosx(1-sinx) / (1-sin^2x) + cos(1+sinx)/(1-sin^2x) now the denominator are the same , we can add the terms

OpenStudy (unklerhaukus):

2cosx/cos^2x is right, now just simplify a little bit

OpenStudy (anonymous):

i dont know how to simplify that

OpenStudy (unklerhaukus):

you have 2cosx/cos^2x if we imagine that cosx = z you have 2z/z^2

OpenStudy (unklerhaukus):

cancel a factor of z (cosine)

OpenStudy (anonymous):

2cosx??

OpenStudy (unklerhaukus):

almost, try again

OpenStudy (anonymous):

2cos?

OpenStudy (unklerhaukus):

no

OpenStudy (anonymous):

sorry its late...real late

OpenStudy (anonymous):

cosx

OpenStudy (unklerhaukus):

2z/z^2 = 2/z

OpenStudy (anonymous):

2/cosx?

OpenStudy (unklerhaukus):

yes, but you can simplify further still

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

sec

OpenStudy (unklerhaukus):

2secx

OpenStudy (anonymous):

the 2 is kept?

OpenStudy (unklerhaukus):

yes,

OpenStudy (anonymous):

cosx=1/secx right?

OpenStudy (unklerhaukus):

that is right

OpenStudy (anonymous):

so its just replacing the 1 with a 2?

OpenStudy (unklerhaukus):

you had 2 / cos x using 1/ cos x = sec x we get 2sec x

OpenStudy (anonymous):

all right one quick question.... if i had cos^4x/sin^2x would it simplify to cot^2x?

OpenStudy (unklerhaukus):

cos^4x / sin^2x = cos^2x * cos^2x / sin^2x = cos^2x * cot^2

OpenStudy (anonymous):

ok thanks a lot!!!